# Daily Land Surface Temperature Reconstruction in Landsat Cross-Track Areas Using Deep Ensemble Learning With Uncertainty Quantification

### v1

@@ V-C Results Under Heavily-Cloudy Situations
The results are shown in Figure 9. Quantitatively, the proposed method achieved LST reconstruction results of R2=0.85-0.98, RMSE=0.84-1.62 K, MAE=0.62-1.23 K, and Bias=-0.67-0.21 K for the three cities. These reconstructions were obtained with cloud coverage of 83.6%-85.6%. Compared to clear-sky situations, the reconstructed LST images under heavily-cloudy situations clearly have some patterns showing the effects of heavy cloud coverage on LST. For example, for the NYC case (Figure 9), the central north areas appear to have relatively lower LST values under heavily-cloudy situations than it is supposed to be under clear-sky situations; and as shown from the image (Figure 9), these areas are most affected by clouds. Similarly, the eastern part of London (Figure 9) and the central part of Hong Kong also shows relatively lower LST values compared to clear-sky situations. These quantitative results and qualitative visualizations demonstrate the success of the proposed method for LST reconstruction under heavily-cloudy situations.
V-D Indirect Validation via Estimating Near-Surface Air Temperature
@@
To further validate the proposed method, we tested the performance of the reconstructed LST in estimating near-surface air temperature. To compare the performance of estimating near-surface air temperature through valid-observed LST (under clear-sky situations) and reconstructed LST (under cloud-covered and no Landsat overpass situations), ideally we can apply them together to build one simple linear model, and the results should be consistent regardless of LST situations. But given that there are more cloud-covered and no Landsat overpass data points, the linear model could be dominated by these data points and thus biased. For validation purposes, we should fit two linear models separately, one for clear-sky valid-observed LST situations, and another for cloud-covered or no Landsat overpass situations. The reconstruction is good if the performance of the latter model has similar capability of the former model. Thus, we built two models using all stations from one city, separately for the two LST situations. We plot the estimated air temperature and in-situ air temperature for each station in Figure 10 for New York City, in Figure 11 for London, and in Figure 12 for Hong Kong. The results, overall and on individual stations, are shown in Tables I,  II,  II, respectively. Overall, the proposed method achieved similar results on using reconstructed LST to estimate near-surface air temperature (R2=0.93, RMSE=2.11 K, MAE=1.72 K), compared to the estimation directly through valid-observed LST (R2=0.93, RMSE=2.02 K, MAE=1.66 K).
V-D1 New York City
@@
On the seven meteorological stations in NYC, the estimation of near-surface air temperature achieved comparable results through valid-observed LST (RMSE=1.42-2.96 K) and reconstructed LST (RMSE=1.65-2.73 K). The performance through reconstructed LST is proportionate to the performance through valid-observed LST, meaning that reconstructed LST has consistent characteristics compared with valid-observed LST. From the scatter plots (Figure 10), we see consistency between the estimated and in-situ near-surface air temperature, except for a few data points at the lower end (below 273.15 K, or 0C), which are likely affected by snow. We should note here that this estimation was through very simple linear regression models and by no means the optimized solution to estimate air temperature. We use this method to compare the quality of LST reconstruction only. In summary, the within 2 K estimation for air temperature demonstrates the success of LST reconstruction and its potential in generating accurate air temperature.
@@
TABLE I: Result of estimating near-surface air temperature from the valid-observed LST situation and from reconstructed LST under cloud-covered or no Landsat overpass situations in New York City. The unit of RMSE and MAE is K.
via Valid-Observed LST via Reconstructed LST
Station R2 RMSE MAE R2 RMSE MAE
KJRB 0.96 2.08 1.64 0.95 2.39 2.08
KEWR 0.96 1.77 1.33 0.96 2.05 1.56
KNYC 0.98 1.98 1.73 0.98 2.07 1.65
KTEB 0.98 1.97 1.69 0.98 1.98 1.58
KJFK 0.97 2.96 2.64 0.96 2.73 2.35
KLDJ 0.97 1.42 1.22 0.97 1.65 1.38
KCDW 0.97 1.65 1.41 0.97 1.74 1.43
NYC 0.93 2.02 1.66 0.93 2.11 1.72
V-D2 London
For London, the fitting result is excellent, regardless of LST situations (Figure 11, Table II). For valid-observed situations, the fitting result combining all stations is with R2 of 0.93, RMSE of 1.63 K, and MAE of 1.38 K. For each station, the range is R2 of 0.95-0.99, RMSE of 1.16-2.05 K, and MAE of 1.0-1.7 K. For cloud-covered or no Landsat overpass situations, the fitting result combining all stations is R2 of 0.93, RMSE of 1.63, and MAE of of 1.37. For each station, the range is R2 of 0.95-0.98, RMSE of 1.06-1.89, and MAE of 0.85-1.58. The performance through reconstructed LST under cloud-covered or no Landsat overpass situations is indistinguishable from the performance through valid-observed LST. In some cases the performance under cloud-cover situations is even slightly better, which can be attributed to the fact that LST under partially-cloudy and overcast conditions has better agreement with air temperature due to the blocking of direct solar radiation [46].
@@ V-D2 London
TABLE II: Result of estimating near-surface air temperature from the valid-observed LST situation and from reconstructed LST under cloud-covered or no Landsat overpass situations in London. The unit of RMSE and MAE is K.
via Valid-Observed LST via Reconstructed LST
R2 RMSE MAE R2 RMSE MAE
EGRB 0.99 1.64 1.38 0.98 1.43 1.18
EGWU 0.96 1.61 1.35 0.96 1.52 1.24
EGLL 0.97 1.16 1.00 0.96 1.18 0.97
EGLC 0.98 1.69 1.53 0.95 1.66 1.39
EGKB 0.95 1.37 1.20 0.97 1.06 0.85
KYL 0.98 2.05 1.70 0.96 1.89 1.58
London 0.93 1.63 1.38 0.93 1.48 1.20
V-D3 Hong Kong
Hong Kong has 23 valid meteorological stations; we built the linear regression models and report results on all stations in Table III, but, for demonstration purposes, only show the scatter plots of the first eight stations (alphabetically) in Figure 12. Overall, the estimation through reconstructed LST (RMSE=1.73 K) is again indistinguishable compared with the estimation through valid-observed LST (RMSE=1.71 K). The successful estimation of near-surface air temperature in all three cities demonstrates the generalization of the proposed method and indirectly validates the proposed method in LST reconstruction, especially under cloudy situations.
@@ V-D3 Hong Kong
TABLE III: Result of estimating near-surface air temperature from the valid-observed LST situation and from reconstructed LST under cloud-covered or no Landsat overpass situations in Hong Kong. The unit of RMSE and MAE is K.
via Valid-Observed LST via Reconstructed LST
Station R2 RMSE MAE R2 RMSE MAE
HKA 0.97 3.15 2.95 0.93 2.99 2.49
HKO 0.95 1.23 1.10 0.92 1.39 1.09
HKP 0.94 1.28 0.96 0.92 1.42 1.13
HPV 0.93 1.38 1.05 0.92 1.62 1.35
JKB 0.93 1.53 1.24 0.93 1.45 1.11
KP 0.95 1.23 0.97 0.92 1.39 1.06
KTG 0.94 1.79 1.53 0.93 1.72 1.36
PEN 0.95 2.45 2.18 0.93 2.26 2.02
SE1 0.93 2.19 1.78 0.94 2.20 1.90
SHA 0.94 2.23 1.68 0.90 2.05 1.48
SKG 0.95 1.46 1.28 0.94 1.40 1.05
SKW 0.95 1.43 1.10 0.94 1.35 1.05
SSP 0.95 1.44 1.17 0.92 1.55 1.16
STY 0.97 0.96 0.72 0.93 1.24 0.95
TC 0.89 1.92 1.43 0.91 1.66 1.20
TMS 0.93 2.07 1.69 0.88 1.68 1.19
TU1 0.95 1.34 1.02 0.92 1.66 1.25
TWN 0.93 1.43 1.10 0.91 1.48 1.05
TY1 0.92 1.49 1.29 0.89 1.60 1.33
VP1 0.90 1.76 1.28 0.91 1.47 1.11
WLP 0.96 1.77 1.59 0.89 1.90 1.52
WTS 0.94 1.39 1.16 0.92 1.44 1.09
YLP 0.92 1.55 1.22 0.89 1.80 1.36
Hong Kong 0.91 1.71 1.34 0.89 1.73 1.32
V-E Uncertainty Quantification
The proposed method, DELAG, is the first of its kind that can offer uncertainty quantification. To demonstrate its effectiveness, we visualize the reconstructed data, the 95% prediction intervals, and the valid-observed Landsat LST values along a vertical line (x-axis=300) in Figure 13. This particular vertical line is selected because it spans a broad range of values, from 286 K to 300 K, presenting a challenging scenario for uncertainty quantification. Within this area, DELAG achieved a RMSE of 0.568 K, with the 95% prediction intervals capturing 92.45% of the test data. Notably, the majority of the Landsat data points (red dots) fall within the 95% prediction intervals. We highlight five clusters in Figure 13 with yellow circle to demonstrate DELAG’s accuracy in estimating the 95% prediction intervals. Effective uncertainty quantification should feature dynamic confidence intervals that are neither too narrow nor too wide: overly narrow intervals suggest the model underestimates the uncertainty, while overly wide intervals indicate ineffective uncertainty quantification with little to none information. In four of the highlighted circles (1st, 2nd, 3rd, and 5th), certain regions exhibit wide 95% prediction intervals, accurately encompassing the observed data. In contrast, the 4th yellow circle and the right peak of the 2nd circle show regions with narrow 95% CI ranges that still successfully capture the observed data. This ability to adaptively adjust the 95% prediction intervals demonstrates the proposed method’s reliability in quantifying uncertainty.
@@ V-E Uncertainty Quantification
Figure 13: Demo showing prediction mean and 95% prediction intervals along a vertical line [x=300] obtained from DELAG. Landsat observed test data are marked as red dots. Yellow circles highlighted successful uncertainty quantification in certain areas with either wide or narrow 95% prediction intervals.
V-F Ablation Analysis
@@
Table IV shows the comparison between the proposed method (DELAG) and baseline methods. The numbers reported here are from the first set of experiments. Compared to the baseline methods, the proposed method DELAG successfully improved the reconstruction, in terms of RMSE from 1.88 K to 0.73 K for NYC, from 3.33 K to 0.96 K for London, and from 2.50 K to 0.74 K for Hong Kong. The addition of GP greatly improved the results by minimizing bias. The addition of ensemble learning improved the model performance slightly and greatly enhanced its capability to quantify uncertainty, thus providing reliable reconstruction.
TABLE IV: Performance comparison between the proposed method and the baseline methods. eATC: 4-parameter enhanced ATC models; eATCe: enhanced ATC with deep ensemble learning; eATC+GP: enhanced ATC plus GP; DELAG (eATCe+GP): enhanced ATC with deep ensemble learning plus GP. The unit of RMSE, MAE and Bias is K.
Metric NYC London Hong Kong
eATC R2 0.93 0.51 0.73
RMSE 1.88 3.33 2.50
MAE 1.55 2.88 2.06
Bias 1.31 2.56 1.92
95% CI Coverage NA NA NA
eATCe R2 0.96 0.84 0.83
RMSE 1.65 2.9 2.27
MAE 1.39 2.75 1.94
Bias 1.31 2.56 1.92
95% CI Coverage 60.53% 57.44% 58.86%
eATC+GP R2 0.93 0.67 0.82
RMSE 1.55 1.86 1.29
MAE 1.19 1.47 1.01
Bias 0.18 0.01 0.18
95% CI Coverage 28.88% 32.18% 31.44%
DELAG R2 0.97 0.9 0.93
RMSE 0.73 0.96 0.74
MAE 0.56 0.74 0.56
Bias -0.01 -0.27 -0.02
95% CI Coverage 97.64% 96.41% 97.95%
VI Conclusion and Discussion
@@
Having daily temperature data at high spatial resolution is critical for many applications. However, existing temperature data, such as gridMET and MODIS, cannot capture the temperature variations within cities. While Landsat is the only satellite series that provides LST data at fine scale (100 m), existing studies have overlooked its potential in generating seamless temperature data. In this study, we leverage the importance of cross-track characteristics and dual-satellite operation of Landsat since 2021 and argue that Landsat is a valuable source for generating seamless temperature data at high spatiotemporal resolution.
@@
We propose DELAG, a deep ensemble learning method incorporating annual temperature cycles (ATC) and Gaussian processes (GP), for Landsat LST reconstruction. We selected three representative cities–New York City, London and Hong Kong–as study areas to show the generalization of the proposed method. We tested the proposed method under three different cloud situations. Under clear-sky situations with real-world cloud patterns, the proposed method achieved LST reconstruction with RMSE of 0.73 K, 0.96 K and 0.74 K for the three cities, respectively. Under heavily-cloudy situations, the proposed method achieved LST reconstruction with RMSE of 0.84 K, 1.62 K and 1.32 K, respectively. These results are superior to existing methods and, additionally, the proposed method can offer uncertainty quantification, providing more reliable reconstructions. We further tested the proposed method’s capability to estimate near-surface air temperature using simple linear regressions. The proposed method successfully estimated near-surface air temperature with RMSE of 2.11 K, 1.48 K and 1.73 K for the three cities respectively through reconstructed LST, comparable to the estimations through valid-observed LST (RMSE of 2.02 K, 1.63 K and 1.71 K). The successful estimations of near-surface air temperature and the comparable results through reconstructed LST and valid-observed LST indirectly validate the proposed method in reconstructing seamless temperature data for real-world applications, such as examining effects of respiratory diseases and wildfire impacts [51, 24]. As existing seamless air temperature products are largely derived from satellite observations, the proposed framework and method thus provides a novel and practical pathway for real-world applications requiring seamless temperature data at high spatiotemporal resolution.
@@
Landsat data frequency plays a major role in LST reconstruction. Landsat Next will greatly enhance data frequency, offering better data availability that is critical for daily LST reconstruction [32]. Enhancing data availability by integrating multiple LST observations can also enhance daily LST reconstruction globally. Satellites and sensors from the Sentinel satellite series, the GOES satellites, and the International Space Station can be of importance. We plan to integrate multiple satellite sources to achieve daily LST reconstruction in the future.

### v2

@@ VI-D Results Under Heavily-Cloudy Situations
The results are shown in Fig. 11. Quantitatively, the proposed method achieved LST reconstruction results of R2=0.85-0.98, RMSE=0.84-1.62 K, MAE=0.62-1.23 K, and Bias=-0.67-0.21 K for the three cities. These reconstructions were obtained with cloud coverage of 83.6%-85.6%. Compared to clear-sky situations, the reconstructed LST images under heavily-cloudy situations clearly have some patterns showing the effects of heavy cloud coverage on LST. For example, for the NYC case (Fig. 11), the central north areas appear to have relatively lower LST values under heavily-cloudy situations than it is supposed to be under clear-sky situations; and as shown from the cloud patterns in Fig. 11, these areas are most affected by clouds. Similarly, the eastern part of London and the central part of Hong Kong (Fig. 11) also show relatively lower LST values compared to clear-sky situations. These quantitative results and qualitative visualizations demonstrate the success of the proposed method for LST reconstruction under heavily-cloudy situations.
VI-E Indirect Validation via Estimating Near-Surface Air Temperature
@@
To further validate the proposed method, we tested the performance of the reconstructed LST in estimating near-surface air temperature. To compare the performance of estimating near-surface air temperature through valid-observed LST (under clear-sky situations) and reconstructed LST (under cloud-covered and no Landsat overpass situations), ideally we can apply them together to build one simple linear model, and the results should be consistent regardless of LST situations. But given that there are more cloud-covered and no Landsat overpass data points, the linear model could be dominated by these data points and thus biased. For validation purposes, we should fit two linear models separately, one for clear-sky or valid-observed LST situations, and another for cloud-covered or no Landsat overpass situations. The reconstruction is good if the performance of the latter model has similar capability of the former model. Thus, we built two models using all stations from one city, separately for the two LST situations. We plot the estimated air temperature and in-situ air temperature for each station in Fig. 12 for New York City, in Fig. 13 for London, and in Fig. 14 for selected stations in Hong Kong. Overall, the proposed method achieved similar results on using reconstructed LST to estimate near-surface air temperature (R2=0.93, RMSE=2.11 K, MAE=1.72 K), compared to the estimation directly through valid-observed LST (R2=0.93, RMSE=2.02 K, MAE=1.66 K).
VI-E1 New York City
@@
On the seven meteorological stations in NYC, the estimation of near-surface air temperature achieved comparable results through valid-observed LST (RMSE=1.42-2.96 K) and reconstructed LST (RMSE=1.65-2.73 K). The performance through reconstructed LST is proportionate to the performance through valid-observed LST, meaning that reconstructed LST has consistent characteristics compared with valid-observed LST. From the scatter plots (Fig. 12), we see consistency between the estimated and in-situ near-surface air temperature, except for a few data points at the lower end (below 273.15 K, or 0C), which are likely affected by snow. We should note here that this estimation was through very simple linear regression models and by no means the optimized solution to estimate air temperature. We use this method to compare the quality of LST reconstruction only. In summary, the within 2 K estimation for air temperature demonstrates the success of LST reconstruction and its potential in generating accurate air temperature.
@@
VI-E2 London
For London, the fitting result is excellent, regardless of LST situations (Fig. 13). For valid-observed situations, the fitting result combining all stations is with R2 of 0.93, RMSE of 1.63 K, and MAE of 1.38 K. For each station, the range is R2 of 0.95-0.99, RMSE of 1.16-2.05 K, and MAE of 1.0-1.7 K. For cloud-covered or no Landsat overpass situations, the fitting result combining all stations is R2 of 0.93, RMSE of 1.63, and MAE of of 1.37. For each station, the range is R2 of 0.95-0.98, RMSE of 1.06-1.89, and MAE of 0.85-1.58. The performance through reconstructed LST under cloud-covered or no Landsat overpass situations is indistinguishable from the performance through valid-observed LST. In some cases the performance under cloud-cover situations is even slightly better, which can be attributed to the fact that LST under partially-cloudy and overcast conditions has better agreement with air temperature due to the blocking of direct solar radiation [69].
@@ VI-E2 London
VI-E3 Hong Kong
We show the scatter plots for the first eight stations (alphabetically) in Fig. 14 (the full result from all 23 stations can be found in the supplementary table). Overall, the estimation through reconstructed LST (RMSE=1.73 K) is again indistinguishable compared with the estimation through valid-observed LST (RMSE=1.71 K). The successful estimation of near-surface air temperature in all three cities demonstrates the generalization of the proposed method and indirectly validates the proposed method in LST reconstruction, especially under cloudy situations.
@@ VI-E3 Hong Kong
Figure 15: Uncertainty Quantification with DELAG along a vertical line at x=300. Landsat observed test data are red dots. Yellow circles highlight reasonable wide or narrow quantified uncertainties.
VI-F Uncertainty Quantification
The proposed method, DELAG, is the first of its kind that can offer uncertainty quantification. To demonstrate its effectiveness, we visualize the reconstructed data, the 95% prediction intervals, and the valid-observed Landsat LST values along a vertical line (x-axis=300) in Fig. 15. This particular vertical line is selected because it spans a broad range of values, from 286 K to 300 K, presenting a challenging scenario for uncertainty quantification. Within this area, DELAG achieved a RMSE of 0.568 K, with the 95% prediction intervals capturing 92.45% of the test data. Notably, the majority of the Landsat data points (red dots) fall within the 95% prediction intervals. We highlight five clusters in Fig. 15 with yellow circles to demonstrate DELAG’s accuracy in estimating the 95% prediction intervals. Effective uncertainty quantification should feature dynamic confidence intervals that are neither too narrow nor too wide: overly narrow intervals suggest the model underestimates the uncertainty, while overly wide intervals indicate ineffective uncertainty quantification with little to none information. In four of the highlighted circles (1st, 2nd, 3rd, and 5th), certain regions exhibit wide 95% prediction intervals, accurately encompassing the observed data. In contrast, the 4th yellow circle and the right peak of the 2nd circle show regions with narrow 95% CI ranges that still successfully capture the observed data. This ability to adaptively adjust the 95% prediction intervals demonstrates the proposed method’s reliability in quantifying uncertainty.
VI-G Comparison with Existing Methods and Ablation Analysis
We compare the proposed DELAG with two existing methods: Deep Feature Gaussian Processes (DFGP) [37] and a 4-step ATC-based model (4-step ATC) [30], along with three baseline models. The results are shown in Table I. The numbers reported here are from the first set of experiments. Across all three datasets, the proposed method (DELAG) consistently outperformed the competitors and the baseline models in almost all metrics.
@@ VI-G Comparison with Existing Methods and Ablation Analysis
Compared to the baseline methods, the proposed method DELAG successfully improved the reconstruction, in terms of RMSE, from 1.88 K to 0.73 K for NYC, from 3.33 K to 0.96 K for London, and from 2.50 K to 0.74 K for Hong Kong. The addition of GP greatly improved the results by minimizing bias. The addition of ensemble learning slightly improved the model performance and greatly enhanced its capability to quantify uncertainty, thus providing reliable reconstruction.
We want to further highlight here that only DELAG is a daily LST reconstruction method. DFGP [37] only works when there are valid observations, and the 4-step ATC model [30] is only effective when there are valid observations within a search window with similar vegetation and land surface. Both models are reduced to ATC only when there are no valid observations. On the contrary, the proposed method, through utilizing ERA5 reanalysis data, yields reasonably good results even under heavily cloudy conditions, as evidence from the previous result on assessing their ability to estimate near-surface air temperature.
TABLE I: Comparison of algorithms on three datasets using MAE, RMSE, $R^{2}$, Bias, and 95% uncertainty coverage (Cov@95). Lower MAE, RMSE, and Bias are better; higher $R^{2}$ and Cov@95 indicate better performance.
Method NYC London Hong Kong
MAE (K) RMSE (K) $R^{2}$ Bias (K) Cov@95 MAE (K) RMSE (K) $R^{2}$ Bias (K) Cov@95 MAE (K) RMSE (K) $R^{2}$ Bias (K) Cov@95
DFGP [37] 1.25 1.69 0.83 -0.24 36.32% 1.02 1.37 0.79 -0.31 39.95% 1.01 1.36 0.75 0.11 38.24%
4-step ATC [30] 0.66 0.90 0.95 -0.40 NA 2.05 2.43 0.61 1.27 NA 0.78 1.07 0.85 -0.48 NA
eATC 1.55 1.88 0.93 1.31 NA 2.88 3.33 0.51 2.56 NA 2.06 2.50 0.73 1.92 NA
eATCe 1.39 1.65 0.96 1.31 60.53% 1.94 2.27 0.83 1.92 58.86% 1.94 2.27 0.83 1.92 58.86%
eATC + GP 1.19 1.55 0.93 0.19 28.88% 1.47 1.86 0.67 0.01 32.18% 1.01 1.29 0.82 0.18 31.44%
DELAG 0.56 0.73 0.97 -0.01 97.64% 0.74 0.96 0.90 -0.27 96.41% 0.56 0.74 0.93 -0.02 97.95%
VI-H Sensitivity of Features in Gaussian Processes
We test the sensitivity of using Landsat and Sentinel-2 mean spectral bands as the features used in Gaussian processes, and the results are shown in Table II. The results from different spectral bands are comparable to each other. Using annual mean spectral bands is the classic option to represent the Earth surface. Recent advances in deep representation learning has shown some promising direction on using features obtained through Earth foundation models for accurate representation [74, 75], which may be integrated here in the future.
TABLE II: Comparison of Spectral Bands (Landsat vs. Sentinel) for Gaussian Process Construction.
Study Area Method MAE (K) RMSE (K) R2
NYC eATCe 1.39 1.65 0.96
DELAG (w/ Landsat) 0.55 0.72 0.97
DELAG (w/ Sentinel) 0.56 0.73 0.97
London eATCe 2.75 2.90 0.84
DELAG (w/ Landsat) 0.67 0.89 0.91
DELAG (w/ Sentinel) 0.74 0.96 0.90
Hong Kong eATCe 1.94 2.27 0.83
DELAG (w/ Landsat) 0.60 0.79 0.92
DELAG (w/ Sentinel) 0.56 0.74 0.93
VI-I Impacts of Landsat Data Frequency
To investigate the impacts of data frequency, we designed a set of experiments with three different observation settings. The three settings are, respectively, dual satellites with Landsat 8 and 9 (4 scenes per 16 days) that is the setting we used in this paper, Landsat 8 with cross-track areas (2 scenes per 16 days) that is a scenario where only one Landsat satellite is available, and Landsat 8 without cross-track areas (1 scene per 16 days). As the numbers of observed samples were not the same across different settings, for fair comparison, we evaluated the performance over their ability to predict near-surface air temperature. For each study area, we combined all data samples (observed or non-observed) together to build one regression model. The results are presented in Table III.
We observe a consistent trend that with more data, the performance in estimating air temperature is better. For NYC and London, with 50% capacity (2 scenes per 16 days), the performance difference is marginal, with RMSE larger by less than 0.1 K. But with 25% capacity (1 scene per 16 days), the performance is significantly worse, degrading to 3.03 K RMSE for NYC and 2.71 K RMSE for London. The performance over Hong Kong is significantly degraded with 50% capacity already, likely due to its more cloudy climate. This set of experiments shows that either cross-track or dual-satellite is important for achieving desirable LST reconstruction using the proposed method.
TABLE III: Impacts of Landsat data frequency.
Setting Scene Frequency NYC London Hong Kong
MAE (K) RMSE (K) $R^{2}$ MAE (K) RMSE (K) $R^{2}$ MAE (K) RMSE (K) $R^{2}$
Landsat 8 + 9, cross-track 4 per 16 days 1.66 2.06 0.94 1.11 1.39 0.94 1.32 1.72 0.89
Landsat 8, cross-track 2 per 16 days 1.74 2.13 0.93 1.18 1.46 0.93 1.81 2.34 0.82
Landsat 8, single-track 1 per 16 days 2.38 3.03 0.86 2.18 2.71 0.76 2.29 2.81 0.77
VII Discussion and Conclusion
@@
Having daily temperature data at high spatial resolution is critical for many applications. Yet, existing temperature data, such as gridMET and MODIS, cannot capture the temperature variations within cities. While Landsat is the only satellite series that provides LST data at fine scale (100 m), existing studies have overlooked its potential in generating seamless temperature data. In this study, we leverage the cross-track characteristics and dual-satellite operation of Landsat since 2021 and argue that Landsat is a valuable source for generating seamless temperature data at high spatiotemporal resolution.
@@
We propose DELAG, a deep ensemble learning method incorporating annual temperature cycles (ATC) and Gaussian processes (GP), for Landsat LST reconstruction. We selected three representative cities—New York City, London and Hong Kong—as study areas to show the generalization of the proposed method. We tested DELAG under three different cloud situations. Under clear-sky situations with real-world cloud patterns, DELAG achieved LST reconstruction with RMSE of 0.73 K, 0.96 K and 0.74 K for the three cities, respectively. Under heavily-cloudy situations, DELAG achieved LST reconstruction with RMSE of 0.84 K, 1.62 K and 1.32 K, respectively. These results are superior to existing methods and, additionally, DELAG can offer uncertainty quantification, providing more reliable reconstructions. We further tested DELAG’s capability to estimate near-surface air temperature using simple linear regressions. The proposed method successfully estimated near-surface air temperature with RMSE of 2.11 K, 1.48 K and 1.73 K for the three cities respectively through reconstructed LST, comparable to the estimations through valid-observed LST (RMSE of 2.02 K, 1.63 K and 1.71 K). The successful estimations of near-surface air temperature and the comparable results through reconstructed LST and valid-observed LST indirectly validate DELAG in reconstructing seamless temperature data for real-world applications, such as examining effects of respiratory diseases and wildfire impacts [16, 76]. As existing seamless air temperature products are largely derived from satellite observations, the proposed framework and method thus provides a novel and practical pathway for real-world applications requiring seamless temperature data at high spatiotemporal resolution.
The LST generated from the proposed method is fundamentally an integration of ERA5 signals and observable pixels (either clear-sky or some pixels observable in cloudy conditions). Overcast conditions are not directly observable, so the associated information is estimated (extrapolated) using daily fluctuations from ERA5-a weather and climate reanalysis product-with a linear multiplier. This approach assumes a linear relationship of the variation from mean (after subtracted by ATC) between local pixels and regional temperature values. While this estimation is reasonable and leverages the best available data, it remains an estimation that may not account for all local weather conditions. Nonetheless, the reconstructed LST has demonstrated far more realistic performance than purely clear-sky LST reconstruction, as evidenced by its performance in estimating near-surface air temperature.
@@
Landsat data frequency plays a major role in LST reconstruction. Landsat Next will greatly enhance data frequency, offering better data availability that is critical for daily LST reconstruction [55]. Enhancing data availability by integrating multiple LST observations can also enhance daily LST reconstruction globally. Satellites and sensors from the Sentinel satellite series, the GOES satellites, and the International Space Station can be of importance. We plan to integrate multiple satellite sources to achieve daily LST reconstruction in the future. Additionally, the reconstructed data here are still LST instead of air temperature. To fully achieve air temperature mapping, a large number of weather stations and a proper cross-validation strategy to separate training and testing stations are necessary.